OpenStudy (anonymous):

What are the real zeros of f(x) = x3 – 3x2 – 4x?Select one of the options below as your answer: A. 0, 1, 4 B. 1, -1, 4 C. -1, 0, 4 D. 1, 0, -4

OpenStudy (anonymous):

when f(x) = 0 x^3 - 3x^2 - 4x = 0 x(x^2 - 3x - 4) = 0 x[(x+1)(x-4)] = 0 So x = -1, 0 or 4 Therefore the answer is C

OpenStudy (anonymous):

thanks!!!!!!

OpenStudy (anonymous):

No problem